Calculate the number of atoms in 0.5 moles atoms of c12.

Grade:11

1 Answers

Ravleen Kaur
5 years ago
Hello

1 mole of c-12 = 6.022 * 10^23 atoms
0.5 mole of c-12 = 3.011*10^23atoms